In the Philippines one of the sources of energy is natural gas like the Malampaya Power Plant. One of the steps in producing electricity from natural gas, pertains to the burning of natural gas to boil water that becomes steam. (a) what chemical reaction takes place ? (b) Word equation: ___+___=___+___ (c) skeleton equation: ___+___=____+___ (d) balance chemical equation: ___+___=___+___(e) how many moles of the reactant and product react with each other ? (f) what are the corresponding mass of each reactant and product ? ___+___=___+___(g) what is being illustrated with the mass of each the substances present ?
1
Expert's answer
2021-10-24T01:25:42-0400

Producing electricity from natural gas, pertains to the burning of natural gas to boil water that becomes steam. Chemical reaction describing this process is:

CH4 + 2O2 = CO2 + 2H2O

Here natural gas burns in oxygen to produce carbon dioxide and water vapor.

1 mole of natural gas requires 2 moles of oxygen to form 1 mole of carbon dioxide and 2 moles of water.

n = m / M

M (CH4) = 16 g/mol

M (O2) = 32 g/mol

M (CO2) = 44 g/mol

M (H2O) = 18 g/mol

If 1 g of CH4 is being burned, the corresponding masses of other reactants are the following:

n (CH4) = 1 / 16 = 0.0625 mol

m = n x M

m (O2) = 0.0625 x 2 x 32 = 4 g

m (CO2) = 0.0625 x 44 = 2.75 g

m (H2O) = 0.0625 x 2 x 18 = 2.25 g

Mass of reactants is equal to the mass of products:

1 + 4 = 2.75 + 2.25